# Retrieval-augmented code completion for local projects using large language models

### v1

@@ Abstract
The use of large language models (LLMs) is becoming increasingly widespread among software developers. However, privacy and computational requirements are problematic with commercial solutions and the use of LLMs. In this work, we focus on using LLMs with around 160 million parameters that are suitable for local execution and augmentation with retrieval from local projects. We train two models based on the transformer architecture, the generative model GPT-2 and the retrieval-adapted RETRO model, on open-source Python files, and empirically evaluate and compare them, confirming the benefits of vector embedding based retrieval. Further, we improve our models’ performance with In-context retrieval-augmented generation, which retrieves code snippets based on the Jaccard similarity of tokens. We evaluate In-context retrieval-augmented generation on larger models and conclude that, despite its simplicity, the approach is more suitable than using the RETRO architecture. We highlight the key role of proper tokenization in achieving the full potential of LLMs in code completion.
@@ 1 Introduction
Programming with the help of large language models (LLMs) and other artificial intelligence (AI) tools is becoming increasingly widespread (Liang et al., 2024). Programmers use AI tools (e.g., GitHub Copilot111https://github.com/features/copilot) to suggest next lines while writing code (Li et al., 2023; Guo et al., 2024; Rozière et al., 2024), to write tests and shorter or longer sections of code (Austin et al., 2021; Lu et al., 2021), and to facilitate learning new technologies (Kasneci et al., 2023). Due to the usefulness of AI tools, code editor developers are rapidly integrating AI into their products.
@@
This raises concerns regarding the privacy of code, which is mostly sent to servers where it is processed by commercial models that can contain hundreds of billions of parameters (Brown et al., 2020). Such large models exceed the computational capabilities available on personal devices. A number of solutions that allow LLMs to run efficiently on local hardware are being developed222An example is available at https://github.com/ggerganov/llama.cpp.. However, the choice of a model size represents an important trade-off, as the model size has a significant impact on the quality of suggestions, the inference speed, and the utilization of compute and memory resources. By focusing on a single programming language or solving just one specific programming task, such as line completion, smaller models can achieve useful results. Additionally, smaller models can speed up the performance of larger models through speculative decoding, where a larger model is used to correct the errors of a smaller model (Leviathan et al., 2023). In our work, we focus on the use of LLMs with fewer parameters to complete lines of Python code. This allows the model to be executed within the code editor on the user’s hardware, even without a GPU to speed up the execution.
@@
LLMs take a string as input and predict its most probable continuation. Usually, the input string or context is the content of the current file being edited. Programming projects are often composed of many files that refer to each other, so we want to enrich the context with information from other files in the project. However, as the context size of models is limited from a few dozen to a hundred lines of code, we need to find informative data to include in the context. In our work, we augment the context of LLMs with relevant code snippets from other files in the project, which we search for based on the similarity of vector embeddings, or alternatively based on the Jaccard similarity of tokens. The goal of our work is to experimentally verify the impact of retrieval-augmented LLMs on the success of completing lines of code. We focus on smaller models suitable for CPU-only inference and on retrieval only from local projects.
@@ 3 Code line completion methods
In this section, we describe our methodology for code line completion based on LLMs. We first describe the retrieval enhanced RETRO approach, followed by our improvements, the first concerning the tokenization, and the second using context-based RAG.
@@ 3.1 The RETRO approach
RETRO (Retrieval-Enhanced Transformer) is a transformer based autoregressive language model enhanced with additional snippets retrieved from a database (Borgeaud et al., 2022). The approach splits the input token sequence into contiguous fixed-size chunks and finds snippets similar to the previous chunk when processing the current chunk. Introducing retrieval into the model allows adapting the model to new data without updating its weights, which is particularly useful for predicting facts not present in the training set (e.g., names in a local programming project). Below, we explain the details of the approach, relevant for our application. Our RETRO approach is illustrated in Figure 1.
@@ 3.2 Token healing
The use of subword tokenization can, in certain cases, lead to undesirable effects during text generation, especially when completing lines of code starting at unnatural word boundaries. Due to the commonly used BPE tokenization algorithm (Sennrich et al., 2016), some substrings appear in multiple tokens. Consequently, if the input string ends with a substring that is the beginning of some token, the output probability distribution for the next token will be biased (Dagan et al., 2024). The tokenization of the string may end with a token that would be merged with the token we are predicting in the BPE algorithm. As a result, tokens that continue the tokenized string have a higher probability than tokens that would better complete the string truncated by the last token. In addition, without a corrective measure, called token healing, may be an out-of-distribution example, leading to unpredictable completion results, since the model has not encountered such unnatural cut-off boundaries during training.
@@ 4.2 The StarCoder dataset
For training our models, we use the StarCoder dataset333 Available at https://huggingface.co/datasets/bigcode/starcoderdata. (Li et al., 2023). The dataset contains over 300 million open-source files from more than 80 programming languages, but we use only files in the Python programming language. We split the dataset into a training and test set in ratio 90%; 10%. We split the data at the project level and not at the file level, so that all files belonging to a project are placed in the same training or test set. The validation set is constructed from 1% of the test set and is used to monitor metrics during model training to detect any potential overfitting (which we did not observe). After training, we evaluate the models’ performance on both the test set and the evaluation sets as described in Section 4.3. Table 2 summarizes the sizes of the training and test sets.
| # Files | # Projects | Total file size | |
@@ 4.3 The RepoEval dataset
For the evaluation of retrieval-augmented line completion on real projects, we use the RepoEval dataset (Zhang et al., 2023b), which contains eight open-source projects from various domains and is divided into three evaluation sets: line, api, and function. The line set contains randomly selected individual lines, while the api set contains lines with API calls to objects defined within the project. These calls can be longer than a single line (10% of examples contain at least 5 lines). The function set is intended for completing function bodies, but we do not use it as it requires completing up to 30 lines and evaluation is performed by executing the projects’ unit tests.
In both cases, line and api, the goal is to predict the entire next line or function call. We want to use our models in a code editor even when the cursor is in the middle of a line, while the user is typing. Therefore, we also evaluate the models’ performance on lineR and apiR, where we use the model to complete code from a random position within the line to the end of the original example.
@@ 4.4 Retrieval database
For building the retrieval database, we use a chunk size of tokens and the CodeT5+ (Wang et al., 2023b) encoder for computing chunk embeddings. To enable faster retrieval of the -nearest chunks, we index the stored embeddings with Faiss (Johnson et al., 2021) as described in C. The pretrained CodeT5+ encoder has 110 million parameters, generates normalized 256-dimensional embeddings and was trained on various code tasks. We also experimented with using the BERT-large encoder with 340 million parameters and 1024-dimensional vector embeddings, as done by Borgeaud et al. (2022) and described in B. However, we opt for the CodeT5+ encoder since it has fewer parameters, outputs smaller dimensional embeddings, and is trained for code understanding, making it a better fit for our use case.
@@ 4.7.1 Single-token metrics
During model training, we utilize the metrics , , and (Zhang et al., 2023a; Jurafsky and Martin, 2009). Let be a sequence of tokens. Perplexity measures the model’s uncertainty about the data and is defined as
@@ 4.7.2 Multi-token predictions
For multi-token predictions, we evaluate the quality of code completion, not only based on the prediction of the next token, but on generating multiple tokens. We calculate metrics on strings, removing leading and trailing whitespace from the strings (Lu et al., 2022; Zhang et al., 2023b). Let be the model’s prediction and the target string. Exact Match is 1 if the strings and match exactly and 0 otherwise. Edit Similarity considers how many edits are needed to transform into :
@@ 5.1 Evaluation on the StarCoder test set
From Table 6, we can see that the single-token metrics using only RETRO with the training set (without the test set) for retrieval are comparable to those of GPT. This can be attributed to the fact that retrieving from the training set does not provide new information, as both models have seen the entire training set during training. We therefore conclude that retrieval is only meaningful on new, previously unseen data for the model. When retrieving from projects, the data is new, as the projects are sampled from the test set. An implication of this observation is that it might be beneficial to train RETRO using a retrieval database different from the training set. The training could also include examples of random chunks to help the model better learn to skip irrelevant information, and examples of same-document snippets to help the model better learn to copy correct answers.
@@ 5.2.3 In-context RAG
Since the found snippets are included directly in the model’s context, it is important that the model supports a sufficiently large context to append the retrieved data in addition to the input context. When using our models with a context size of 384 tokens, we often need to truncate the input sequence to leave enough space to include the retrieved snippets at the beginning of the context. For our models, GPT-2 and RETRO, we include only one retrieved snippet due to the small context size of the model. For models with larger context sizes, we include two retrieved snippets. In Table 7 and in Figures 6 and 7, we report multi-token prediction metrics both with and without retrieval as described in Section 3.3. The results indicate that using In-context RAG leads to improvements in multi-token predictions compared to the baseline models in all cases.
@@
In Section 5.1, we estimate the upper bound of single-token metrics for RETRO by not removing chunks retrieved from the same file as the input context. Similarly, we want to determine the upper bound of model performance when using In-context RAG if the retrieved snippets can contain the continuation of the current file we are completing. This means that the context includes the line we want to complete, allowing the model to simply copy it from the context. In Figure 8, we see that copying from retrieved snippets is very effective, while copying with RETRO without In-context RAG performs similarly to using RETRO with In-context RAG without copying. In addition to the explanations in Section 5.1, we also note that because the first Retro layer is the sixth layer (which is also the last and only Retro layer for RETRO), certain information from the input tokens may get diluted, preventing direct copying of tokens from the encoded chunks with chunked cross-attention. We hypothesize that adding copying examples to the training would improve the model’s ability to copy, and Retro layers could be included closer to the beginning of processing instead of only in the sixth layer. Accurate copying is important because when completing code, we want to extract not only the semantic meaning from the found snippets but also exact facts (e.g., the exact name of an existing function and not a hallucinated but semantically similar name).
@@ 6 Conclusion
A major advantage of In-context RAG is that it does not require updating the model weights, but doing so could still be beneficial for improving performance either by fine-tuning the model with included retrieved snippets or by fine-tuning the retriever (Karpukhin et al., 2020; Izacard and Grave, 2021). For more efficient execution in code editors, models could be fine-tuned to append to the end of the context instead of the beginning, which would improve the caching of keys and values in self-attention during frequent code editing. A next step would also be to explore the impact of combining In-context RAG with the fill-in-the-middle technique, which is becoming increasingly widespread (Wu et al., 2024).
@@ Acknowledgements
Marko Hostnik: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Writing – original draft. Marko Robnik-Šikonja: Conceptualization, Writing – original draft, Writing – review and editing, Supervision, Project administration.
@@
This research was funded by the Slovenian Research and Innovation Agency (ARIS) core research programme P6-0411 and projects J7-3159, CRP V5-2297, L2-50070, and PoVeJMo. (Marko Robnik-Šikonja).
@@ Appendix A StarCoder dataset analysis
Appendix B Chunk embeddings with BERT
@@
Appendix C Indexing chunk embeddings
Due to the large number of chunks in the database, exhaustive searching for the exact -nearest chunks is not feasible, so we resort to approximate searching instead. As a consequence of approximate search, we seek a compromise between the search accuracy, search speed, index building speed, and index memory usage. In our implementation, we use the Faiss library (Johnson et al., 2021) to index the chunk embeddings444In Faiss, the used index can be compactly described by the string OPQ32_128,IVF1048576_HNSW32,PQ32.. The search is sped up using IVF indexing (Johnson et al., 2021) in combination with the HNSW (Malkov and Yashunin, 2020) algorithm. Additionally, the vectors are compressed with the PQ (Jegou et al., 2008) vector codec and the OPQ (Ge et al., 2013) transformation. Table 10 summarizes the space requirements for storing the database and retrieval index based on the tokenization method used.

### v2

@@ Abstract
The use of large language models (LLMs) is becoming increasingly widespread among software developers. However, privacy and computational requirements are problematic with commercial solutions and the use of LLMs. In this work, we focus on using relatively small and efficient LLMs with 160M parameters that are suitable for local execution and augmentation with retrieval from local projects. We train two open transformer-based models, the generative GPT-2 and the retrieval-adapted RETRO, on open-source Python files, and empirically compare them, confirming the benefits of embedding-based retrieval. Furthermore, we improve our models’ performance with In-context retrieval-augmented generation (RAG), which retrieves code snippets using the Jaccard similarity of tokens. We evaluate In-context RAG on larger models and determine that, despite its simplicity, the approach is more suitable than using the RETRO architecture. Experimental results indicate that In-context RAG improves the code completion baseline by over 26%, while RETRO improves over the similarly sized GPT-2 baseline by 12%. We highlight the key role of proper tokenization in achieving the full potential of LLMs in code completion.
@@ 1 Introduction
Programming with the help of large language models (LLMs) and other artificial intelligence (AI) tools is becoming increasingly widespread (Liang et al., 2024). Programmers use AI tools (e.g., GitHub Copilot333https://github.com/features/copilot) to suggest next lines while writing code (Li et al., 2023; Guo et al., 2024; Rozière et al., 2024), to write tests and shorter or longer sections of code (Austin et al., 2021; Lu et al., 2021), and to facilitate learning new technologies (Kasneci et al., 2023).
As AI tools prove increasingly useful, code editor developers are rapidly incorporating AI into their products. However, the widespread adoption of AI tools raises concerns about code privacy, as code is transmitted to remote servers for processing by commercial models that can contain hundreds of billions of parameters (Brown et al., 2020). These large models require computational resources far beyond what ordinary consumer hardware provides, which is why a number of solutions that allow LLMs to run efficiently on consumer hardware are being developed.444An example is available at https://github.com/ggerganov/llama.cpp.
However, the choice of a model size represents an important trade-off, as the model size has a significant impact on the quality of suggestions, the inference speed, and the utilization of compute and memory resources. While existing approaches have focused on improving suggestions using larger models augmented with retrieval from large retrieval datasets (Borgeaud et al., 2022), our approach differs in using small models augmented with only locally available project files.
@@
By focusing on a single programming language or solving just one specific programming task, such as line completion, smaller models can achieve useful results. Additionally, smaller models can speed up the performance of larger models through speculative decoding, where a larger model is used to correct the errors of a smaller model (Leviathan et al., 2023). In our work, we focus on the use of LLMs with fewer parameters to complete lines of Python code. This allows the model to be executed within the code editor on the user’s hardware, even without a graphics processing unit (GPU) to speed up the execution.
@@
LLMs take a string as input and predict its most probable continuation. Usually, the input string or context is the content of the current file being edited. Programming projects are often composed of many files that refer to each other, so we want to enrich the context with information from other files in the project. However, as the context size of models is limited from a few dozen to a hundred lines of code, we need to find informative data to include in the context. In our work, we augment the context of LLMs with relevant code snippets from other files in the project, which we search for based on the similarity of vector embeddings, or alternatively based on the Jaccard similarity of tokens. The goal of our work is to experimentally verify the impact of retrieval-augmented LLMs on the success of completing lines of code. We focus on smaller models suitable for inference without a GPU and on retrieval only from local projects.
Our proposed approach has widespread use cases in improving the coding experience for developers. Small coding LLMs can be used to offer intelligent code completion suggestions in various environments, such as in local text editors. In our work, we demonstrate how small LLMs can be improved and made more viable with retrieval-augmented generation (RAG) using a simple Jaccard similarity-based approach that doesn’t require expensive indexing for retrieving relevant code snippets from developers’ projects. Better code suggestions with smaller LLMs can not only help developers write code faster, but our approach does so while maintaining a focus on privacy without incurring additional costs of commercial LLMs.
@@ 3 Code line completion methods
In this section, we describe our methodology for code line completion based on LLMs. We first describe the retrieval-enhanced RETRO approach, followed by our improvements, the first concerning the tokenization, and the second using context-based retrieval-augmented generation.
@@ 3.1 The RETRO approach
RETRO (Retrieval-Enhanced Transformer) is a transformer based autoregressive language model enhanced with additional snippets retrieved from a database (Borgeaud et al., 2022). The approach splits the input token sequence into contiguous fixed-size chunks and finds snippets similar to the previous chunk when processing the current chunk. Introducing retrieval into the model allows adapting the model to new data without updating its weights, which is particularly useful for predicting facts not present in the training set (e.g., names in a local programming project). For our experiments, we make minor modifications to the original RETRO architecture. We change the chunk embedding model from BERT (Devlin et al., 2019) to CodeT5+ (Wang et al., 2023b) and modify the hyperparameters described in Section 4.5. During inference, we modify the input context by adding some padding tokens, using token healing (see Section 3.2) and In-context RAG (see Section 3.3). Our RETRO approach is illustrated in Figure 1. Below, we explain the details of the approach, relevant for our application.
@@ 3.2 Token healing
The use of subword tokenization can, in certain cases, lead to undesirable effects during text generation, especially when completing lines of code starting at unnatural word boundaries. Due to the commonly used byte pair encoding (BPE) tokenization algorithm (Sennrich et al., 2016), some substrings appear in multiple tokens. Consequently, if the input string ends with a substring that is the beginning of some token, the output probability distribution for the next token will be biased (Dagan et al., 2024). The tokenization of the string may end with a token that would be merged with the token we are predicting in the BPE algorithm. As a result, tokens that continue the tokenized string have a higher probability than tokens that would better complete the string truncated by the last token. In addition, without a corrective measure, called token healing, may be an out-of-distribution example, leading to unpredictable completion results, since the model has not encountered such unnatural cut-off boundaries during training.
@@ 4.2 The StarCoder dataset
For training our models, we use the StarCoder dataset555 Available at https://huggingface.co/datasets/bigcode/starcoderdata. (Li et al., 2023). The dataset contains over 300 million open-source files from more than 80 programming languages, but we use only files in the Python programming language. We split the dataset into a training and test set in ratio 90%; 10%. We split the data at the project level and not at the file level, so that all files belonging to a project are placed in the same training or test set. The validation set is constructed from 1% of the test set and is used to monitor metrics during model training to detect any potential overfitting (which we did not observe). After training, we evaluate the models’ performance on both the test set and the evaluation sets as described in Section 4.3. Table 2 summarizes the sizes of the training and test sets.
| Split | # Files | # Projects | Total file size |
@@ 4.3 The RepoEval dataset
For the evaluation of retrieval-augmented line completion on real projects, we use the RepoEval dataset (Zhang et al., 2023b), which contains eight open-source projects from various domains and is divided into three evaluation sets: line, api, and function. The line set contains randomly selected individual lines, while the api set contains lines with application programming interface (API) calls to objects defined within the project. These calls can be longer than a single line (10% of examples contain at least 5 lines). The function set is intended for completing function bodies, but we do not use it as it requires completing up to 30 lines and evaluation is performed by executing the projects’ unit tests.
In both cases, line and api, the goal is to predict the entire next line or function call. We want to use our models in a code editor even when the cursor is in the middle of a line, while the user is typing. Therefore, we also evaluate the models’ performance on lineR and apiR, where we use the model to complete code from a random position within the line to the end of the original example. More details on lineR and apiR can be found in B and an example in Figure 4.
@@ 4.4 Retrieval database
For building the retrieval database, we use a chunk size of tokens and the CodeT5+ (Wang et al., 2023b) encoder for computing chunk embeddings. To enable faster retrieval of the -nearest chunks, we index the stored embeddings with Faiss (Johnson et al., 2021) as described in D. The pretrained CodeT5+ encoder has 110 million parameters, generates normalized 256-dimensional embeddings and was trained on various code tasks. We also experimented with using the BERT-large encoder with 340 million parameters and 1024-dimensional vector embeddings, as done by Borgeaud et al. (2022) and described in C. However, we opt for the CodeT5+ encoder since it has fewer parameters, outputs smaller dimensional embeddings, and is trained for code understanding, making it a better fit for our use case.
@@ 4.7.1 Single-token metrics
During model training, we use perplexity (), recall-at-k () and mean reciprocal rank () metrics (Zhang et al., 2023a; Jurafsky and Martin, 2009). While is a widespread metric, we find that recall-based metrics provide more intuitive interpretations of results compared to .
Let be a sequence of tokens. Perplexity measures the model’s uncertainty about the data and is defined as
@@ 4.7.2 Multi-token predictions
For multi-token predictions, we evaluate the quality of code completion, not only based on the prediction of the next token, but on generating multiple tokens. We calculate metrics on strings, removing leading and trailing whitespace from the strings (Lu et al., 2022; Zhang et al., 2023b).
Let be the model’s prediction and the target string. Exact Match is 1 if the strings and match exactly and 0 otherwise. Edit Similarity considers how many edits are needed to transform into :
@@ 5.1 Evaluation on the StarCoder test set
From Table 6, we see that the single-token metrics using only RETRO with the training set (without the test set) for retrieval are comparable to those of GPT. This can be attributed to the fact that retrieving from the training set does not provide new information, as both models have seen the entire training set during training. We therefore conclude that retrieval is only meaningful on new, previously unseen data for the model. When retrieving from projects, the data is new, as the projects are sampled from the test set. An implication of this observation is that it might be beneficial to train RETRO using a retrieval database different from the training set. The training could also include examples of random chunks to help the model better learn to skip irrelevant information, and examples of same-document snippets to help the model better learn to copy correct answers.
@@ 5.2.3 In-context RAG
Since the found snippets are included directly in the model’s context, it is important that the model supports a sufficiently large context to append the retrieved data in addition to the input context. When using our models with a context size of 384 tokens, we often need to truncate the input sequence to leave enough space to include the retrieved snippets at the beginning of the context. For our models, GPT-2 and RETRO, we include only one retrieved snippet due to the small context size of the model. For models with larger context sizes, we include two retrieved snippets. In Table 7 and in Figures 7 and 8, we report multi-token prediction metrics both with and without retrieval as described in Section 3.3. The results indicate that using In-context RAG leads to improvements in multi-token predictions compared to the baseline models in all cases.
@@
In Section 5.1, we estimate the upper bound of single-token metrics for RETRO by not removing chunks retrieved from the same file as the input context. Similarly, we want to determine the upper bound of model performance when using In-context RAG if the retrieved snippets can contain the continuation of the current file we are completing. This means that the context includes the line we want to complete, allowing the model to simply copy it from the context. In Figure 9, we see that copying from retrieved snippets is very effective, while copying with RETRO without In-context RAG performs similarly to using RETRO with In-context RAG without copying.
In addition to the explanations in Section 5.1, we also note that because the first Retro layer is the sixth layer (which is also the last and only Retro layer for RETRO), certain information from the input tokens may get diluted, preventing direct copying of tokens from the encoded chunks with chunked cross-attention. We hypothesize that adding copying examples to the training would improve the model’s ability to copy, and Retro layers could be included closer to the beginning of processing instead of only in the sixth layer. Accurate copying is important because when completing code, we want to extract not only the semantic meaning from the found snippets but also exact facts (e.g., the exact name of an existing function and not a hallucinated but semantically similar name).
@@ 6 Conclusion and future work
In summary, users looking to apply our work would do well to consider picking the largest model size that fits within their constraints and then apply retrieval from local projects.
Future work
@@
A major advantage of In-context RAG is that it does not require updating the model weights, but doing so could still be beneficial for improving performance either by fine-tuning the model with included retrieved snippets or by fine-tuning the retriever (Karpukhin et al., 2020; Izacard and Grave, 2021). Instead of retrieving snippets based on the Jaccard similarity of tokens, a future area to investigate is incorporating sparse search with the benefits of dense semantic search and check whether hybrid generative and retrieval-based models could further improve performance. For more efficient execution in code editors, models could be fine-tuned to append to the end of the context instead of the beginning, which would improve the caching of keys and values in self-attention during frequent code editing. A next step would also be to explore the impact of combining In-context RAG with the fill-in-the-middle technique, which is becoming increasingly widespread (Wu et al., 2024).
@@ Acknowledgements
Marko Hostnik: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Writing – original draft, Writing – review and editing. Marko Robnik-Šikonja: Conceptualization, Writing – original draft, Writing – review and editing, Supervision, Project administration.
@@
This research was funded by the Slovenian Research and Innovation Agency (ARIS) core research programme P6-0411 and project GC-0002. The work was also supported by the EU through ERA Chair grant no. 101186647 (AI4DH) and cofinancing for research innovation projects in support of green transition and digitalisation (project PoVeJMo, no. C3.K8.IB) (Marko Robnik-Šikonja).
@@ Appendix A StarCoder dataset analysis
Appendix B RepoEval modification details
The line and api datasets from RepoEval (Zhang et al., 2023b) each contain 1600 examples. Every example in line is composed of a prompt input string and a ground_truth output string. Our modified dataset lineR is constructed from line by expanding each original example’s prompt with a random number of characters from ground_truth and shrinking ground_truth accordingly. We are careful that at least one non-whitespace character from ground_truth is included into the new prompt to avoid expanding prompt only by the leading indentation whitespace. The construction of apiR from api is similar to that of lineR from line.
Appendix C Chunk embeddings with BERT
@@
Appendix D Indexing chunk embeddings
Due to the large number of chunks in the database, exhaustive searching for the exact -nearest chunks is not feasible, so we resort to approximate searching instead. As a consequence of approximate search, we seek a compromise between the search accuracy, search speed, index building speed, and index memory usage. In our implementation, we use the Faiss library (Johnson et al., 2021) to index the chunk embeddings.666In Faiss, the used index can be compactly described by the string OPQ32_128,IVF1048576_HNSW32,PQ32. The search is sped up using IVF indexing (Johnson et al., 2021) in combination with the HNSW (Malkov and Yashunin, 2020) algorithm. Additionally, the vectors are compressed with the PQ (Jegou et al., 2008) vector codec and the OPQ (Ge et al., 2013) transformation. Table 10 summarizes the space requirements for storing the database and retrieval index based on the tokenization method used.